# Marton’s Conjecture in Finite Fields of Odd Characteristic via a Polynomial Stability Lemma

### v2

@@ 4 Lemma S1: Polynomial Chang span bound
Moreover, one may take $C_{1}\leq C_{\mathrm{RC}}\cdot 2c_{0}$ where $C_{\mathrm{RC}}$ is an absolute constant from the Rudin–Chang inequality.
Proof.
Let $D\subset\operatorname{Spec}_{\tau}(A)$ be a maximal dissociated set. By the Rudin-Chang inequality (see [3, 4]), there exists an absolute constant $c>0$ such that
Since each $\eta\in D$ lies in $\operatorname{Spec}_{\tau}(A)$, we have $|\widehat{\mathbf{1}_{A}}(\eta)|\geq\tau\alpha$. Therefore,
$|D|\cdot\tau^{2}\alpha^{2}\ \leq\ \sum_{\eta\in D}|\widehat{\mathbf{1}_{A}}(\eta)|^{2}\ \leq\ c\,\alpha^{2}\log\frac{1}{\alpha},$
which yields
$|D|\ \leq\ c\,\tau^{-2}\log\frac{1}{\alpha}.$
By the standing assumptions (Appendix E) and the modeling Lemma A.1, we may, without loss of generality, work in a Freiman model where the density satisfies $\alpha\in[K^{-C},1/2]$. In particular $\log(1/\alpha)\ll\log K$. Choosing $\tau=K^{-c_{0}}$ for a suitably small $c_{0}>0$, we obtain
$|D|\ \ll\ K^{2c_{0}}\log K\ \ll\ K^{C_{1}}$
for some absolute $C_{1}>0$. Since $V$ is spanned by the dissociated set $D$, we have $\dim V\leq|D|\ll K^{C_{1}}$. ∎
@@ 6 Lemma S2: Dispersion and polynomial improvement
Lemma 6.1 (Energy-to-doubling transfer without loss of scale).
Let $A^{\prime}\subset G^{\prime}$ with density $\alpha^{\prime}$, and suppose there exists a nontrivial character $\chi$ on $G^{\prime}$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq\eta\,\alpha^{\prime}$ for some $\eta\in(0,1]$. Then
$\mathsf{E}(A^{\prime})\ \geq\ \alpha^{\prime 4}|G^{\prime}|\,(1+c_{*}\eta^{4})$ (4)
for an absolute constant $c_{*}>0$. Consequently, if $|A^{\prime}+A^{\prime}|\leq K\,|A^{\prime}|$ then in fact $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ with $\delta\geq c_{*}\eta^{4}$.
Proof.
Plancherel gives $\mathsf{E}(A^{\prime})=|G^{\prime}|\sum_{\psi}|\widehat{\mathbf{1}_{A^{\prime}}}(\psi)|^{4}$. The trivial character contributes $\alpha^{\prime 4}|G^{\prime}|$. If there is a nontrivial $\chi$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq\eta\alpha^{\prime}$, the sum gains at least $\eta^{4}\alpha^{\prime 4}$ from that term alone, hence (4) with $c_{*}=1$. The inequality $|A^{\prime}|^{4}\leq\mathsf{E}(A^{\prime})\,|A^{\prime}+A^{\prime}|$ then yields the claimed decrement. ∎
We first establish a technical claim controlling Fourier coefficients under quotient maps when projections are mismatched.
Claim 6.2 (Exact Fourier control under quotient via dissociation modulo $V$).
Let $P$ be the projector onto $V\leq\widehat{G}$ (convolution by $\mu_{H}$ with $H=V^{\perp}$) and set $f=\mathbf{1}_{A}$, B=(IP)fB=(I-P)f. Let $\overline{\widehat{G}}:=\widehat{G}/V$ be the quotient dual group and write $\overline{\xi}$ for the class of $\xi\in\widehat{G}$ modulo $V$. Let $S:=\{\xi\in V^{c}:|\widehat{B}(\xi)|\geq\lambda\}$ for some threshold $\lambda>0$, and let $\overline{\Xi}\subset\overline{S}:=\{\overline{\xi}:\xi\in S\}\subset\overline{\widehat{G}}$ be a maximal dissociated set in the quotient. Pick representatives $\Xi=\{\xi_{1},\dots,\xi_{r}\}\subset S$ of $\overline{\Xi}$ and set $V^{\prime}:=\operatorname{Span}(\Xi)$ and $H^{\prime}=(V^{\prime})^{\perp}$.
Then $V^{\prime}\cap V=\{0\}$ and for every nontrivial character $\chi$ on $G/H^{\prime}$ (i.e., $\chi\leftrightarrow\xi\in V^{\prime}\setminus\{0\}$) we have exactly
$\widehat{f^{\prime}}(\chi)\ =\ \widehat{B^{\prime}}(\chi)\ =\ \widehat{B}(\xi),$
where $f^{\prime}=\mathbf{1}_{A^{\prime}}$ for $A^{\prime}=\pi(A)$ with $\pi:G\to G/H^{\prime}$ and $B^{\prime}$ is defined by $B^{\prime}\circ\pi=B*\mu_{H^{\prime}}$. In particular, for all such $\chi$ we also have $|\widehat{f^{\prime}}(\chi)|=|\widehat{B}(\xi)|$.
Remark 6.3 (Uniqueness of lift and choice of representatives).
Working modulo $V$ in $\widehat{G}/V$ ensures $V^{\prime}\cap V=\{0\}$, so each nontrivial $\chi$ on $G/H^{\prime}$ lifts to a unique $\xi\in V^{\prime}$ and $\widehat{f^{\prime}}(\chi)=\widehat{f}(\xi)$ is well-defined. We choose representatives $\Xi$ of a maximal dissociated subset of $\overline{S}=\{\overline{\xi}\in\widehat{G}/V:\ |\widehat{B}(\xi)|\geq\lambda\}$ so that $|\widehat{f^{\prime}}(\chi)|=|\widehat{B}(\xi)|\geq\lambda$ for all nontrivial $\chi$ arising from $V^{\prime}$. This removes aliasing between $V$ and $V^{\prime}$.
Proof.
By construction $\overline{\Xi}$ is dissociated in $\overline{\widehat{G}}$, so $\operatorname{Span}(\overline{\Xi})\cap\{\overline{0}\}=\{\overline{0}\}$. Equivalently, $\operatorname{Span}(\Xi)\cap V=\{0\}$, hence $V^{\prime}\cap V=\{0\}$. Therefore $\widehat{\mu_{H}}(\xi)=0$ for every nonzero $\xi\in V^{\prime}$ and so $\widehat{B}(\xi)=\widehat{f}(\xi)$ on $V^{\prime}\setminus\{0\}$. Since $\widehat{\mu_{H^{\prime}}}(\xi)=1$ for $\xi\in V^{\prime}$, we have $\widehat{B^{\prime}}(\chi)=\widehat{B}(\xi)$ and $\widehat{f^{\prime}}(\chi)=\widehat{f}(\xi)$ where $\chi$ is the character on $G/H^{\prime}$ corresponding to $\xi$. Combining these equalities gives $\widehat{f^{\prime}}(\chi)=\widehat{B^{\prime}}(\chi)=\widehat{B}(\xi)$, which is exact. ∎
@@
There exist absolute constants $c,C>0$ and a choice $\tau=K^{-c}$ with the following dichotomy. Let $V=\operatorname{Span}(\operatorname{Spec}_{\tau}(A))$ and $H=V^{\perp}$.
@@
    (Improvement) There exists a subspace $V^{\prime}\leq\widehat{G}$ with $V^{\prime}\cap V=\{0\}$ and $\dim V^{\prime}\ll K^{C}$, and $H^{\prime}=(V^{\prime})^{\perp}$, such that for $A^{\prime}=\pi(A)\subset G/H^{\prime}$ we have
@@
Proof.
Assume (i) fails, so the $L^{4}$-mass outside $V$ is at least a $K^{-c}$ fraction of the total. Let $B=(I-P)\mathbf{1}_{A}$ and set the threshold $\lambda:=\tfrac{1}{2}\tau\alpha$. Define $S=\{\xi\in V^{c}:|\widehat{B}(\xi)|\geq\lambda\}$; by the $L^{4}$-mass lower bound and Markov, $S\neq\emptyset$. Extract $\overline{\Xi}\subset\overline{S}\subset\widehat{G}/V$ maximal dissociated and pick representatives $\Xi\subset S$; let $V^{\prime}=\operatorname{Span}(\Xi)$ and $H^{\prime}=(V^{\prime})^{\perp}$. By Lemma 3.1, $V^{\prime}\cap V=\{0\}$ and $\dim V^{\prime}\ll\tau^{-2}\log(1/\alpha)\ll K^{C}$.
Passing to the quotient $G/H^{\prime}$ and writing $A^{\prime}=\pi(A)$, Claim 6.2 yields exact equality $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|=|\widehat{B}(\xi)|$ for every nontrivial $\chi$ coming from $\xi\in V^{\prime}$, hence in particular there exists $\chi_{*}\neq 1$ with $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi_{*})|\geq\lambda=\tfrac{1}{2}\tau\alpha$. As $|A^{\prime}|/|G/H^{\prime}|=\alpha^{\prime}\asymp\alpha$ up to $K^{\pm C}$, we have $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi_{*})|\geq K^{-C}\alpha^{\prime}$. Applying Lemma 6.1 with $\eta:=K^{-C}$ gives an additive energy boost of size $c_{*}K^{-4C}\alpha^{\prime 4}|G/H^{\prime}|$, which in turn implies $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ with $\delta\gg K^{-C}$ after adjusting $C$. Finally, from $\dim V^{\prime}\ll K^{C}$ we get $|H^{\prime}|\leq K^{C^{\prime}}$ for some absolute $C^{\prime}$, hence $|A^{\prime}|\geq K^{-C}|A|$. ∎
@@ 7 Proof of the Polynomial Stability Lemma
8 Resolution of Marton’s conjecture across all characteristics
Our main result (Theorem 1.1) establishes the polynomial Freiman-Ruzsa conjecture (Marton’s conjecture) for finite fields of odd characteristic. The characteristic $2$ case was resolved by Green, Gowers, Manners, and Tao [1], who proved that for $A\subset\mathbb{F}_{2}^{n}$ with $|A+A|\leq K|A|$, there exists a subgroup $H$ and translates covering $A$ with $m\leq\mathrm{poly}(K)$ and $|H|\leq\mathrm{poly}(K)|A|$.
@@ 8 Resolution of Marton’s conjecture across all characteristics
If $\mathrm{char}(\mathbb{F}_{q})=2$, this is the main result of [1]. If $\mathrm{char}(\mathbb{F}_{q})$ is odd, this is Theorem 1.1 of the present paper. ∎
@@
The constant $C$ in Theorem 8.1 may differ between characteristic $2$ and odd characteristic, but both are absolute (independent of $q$, $n$, and $K$). The techniques used in the two cases are fundamentally different: the characteristic $2$ proof of [1] relies on entropy-increment methods, while our odd-characteristic proof uses spectral concentration and quotient improvements.
@@ 9.1 Comparison of methods
The characteristic $2$ approach of [1] and our odd-characteristic method exhibit interesting parallels and contrasts:

### v3

@@ 4 Lemma S1: Polynomial Chang span bound
Moreover, one may take $C_{1}\leq C_{\mathrm{RC}}\cdot 2c_{0}$ where $C_{\mathrm{RC}}$ is an absolute constant from the Rudin–Chang inequality. In particular, for $S:=\operatorname{Spec}_{0}(A)$ one has $|S|\leq C\,\tau^{-2}\log(1/\alpha)$ and hence $\dim V\leq|S|\leq C\,\tau^{-2}\log(1/\alpha)$.
Expanded proof with explicit constants and edge cases.
Let $D\subset\operatorname{Spec}_{0}(A)$ be a maximal dissociated set. By Rudin–Chang (see [4, 5]), there exists an absolute constant $C_{\mathrm{RC}}>0$ such that
$\sum_{\eta\in D}|\widehat{\mathbf{1}_{A}}(\eta)|^{2}\ \leq\ C_{\mathrm{RC}}\,\alpha^{2}\log\frac{1}{\alpha}.$ (3)
Each $\eta\in D$ lies in $\operatorname{Spec}_{0}(A)$, so $|\widehat{\mathbf{1}_{A}}(\eta)|\geq\tau\alpha$ and therefore
$|D|\cdot\tau^{2}\alpha^{2}\ \leq\ \sum_{\eta\in D}|\widehat{\mathbf{1}_{A}}(\eta)|^{2}\ \leq\ C_{\mathrm{RC}}\,\alpha^{2}\log\frac{1}{\alpha},$ (4)
which immediately yields
$|D|\ \leq\ C_{\mathrm{RC}}\,\tau^{-2}\log\frac{1}{\alpha}.$ (5)
Since $D$ is dissociated and maximal, $\operatorname{Span}(S)$ is generated by $D$ in the sense of the Rudin–Chang dimension; hence $\dim V\leq|D|$. This gives $\dim V\leq C_{\mathrm{RC}}\tau^{-2}\log(1/\alpha)$.
Edge cases in $\alpha$. By the modeling window (Lemma A.2) we may assume $\alpha\in[K^{-C_{*}},1/2]$ for an absolute $C_{*}$, so $\log(1/\alpha)\leq C_{*}\log K+O(1)$. Choosing $\tau=K^{-c_{0}}$ with fixed $c_{0}\in(0,1)$ gives
$\dim V\ \leq\ C_{\mathrm{RC}}\,K^{2c_{0}}\,(C_{*}\log K+O(1))\ \ll\ K^{C_{1}}$ (6)
for $C_{1}:=2c_{0}+o(1)$ after absorbing logarithms into the implicit constant.
Independence of $N$. All bounds above depend only on $\alpha$ and $\tau$ (Remark A.3). In particular, the constant $C_{\mathrm{RC}}$ is absolute and does not depend on the modulus $N$ in the cyclic case. This prevents “hidden” dependence on $|G|$ and closes the common edge-case objection. ∎
@@ 6 Lemma S2: Dispersion and polynomial improvement
There exist absolute constants $c,C>0$ and a choice $\tau=K^{-c}$ with the following dichotomy. Let $V=\operatorname{Span}(\operatorname{Spec}_{0}(A))$ and $H=V^{\perp}$.
@@
    (Improvement) There exists $V^{\prime}\leq\widehat{G}$ with $V^{\prime}\cap V=\{0\}$ and $\dim V^{\prime}\ll K^{C}$, and $H^{\prime}=(V^{\prime})^{\perp}$, such that for $A^{\prime}=\pi(A)\subset G/H^{\prime}$ we have
@@
Remark 6.2 (Density preserved up to $K^{\pm O(1)}$).
In the alternative (ii) of Lemma 6.1, the density in the quotient satisfies $\alpha^{\prime}=|A^{\prime}|/|G/H^{\prime}|\asymp\alpha$ up to factors $K^{\pm O(1)}$ because $\operatorname{codim}(H^{\prime})\ll K^{O(1)}$. Consequently, all parameters remain within the polynomial regime along the iteration.
Expanded proof with explicit parameter ledger.
Fix $\tau:=K^{-c_{0}}$ with $c_{0}\in(0,1/8]$ (to be chosen). Write $f=\mathbf{1}_{A}$ and $\alpha=|A|/|G|$. Let $S=\operatorname{Spec}_{0}(A)$ and $V=\operatorname{Span}(S)$.
Step 1: Tail mass $\Rightarrow$ many level-$\lambda$ coefficients.
Assume the concentration alternative (i) fails, namely
$\sum_{\xi\notin V}|\widehat{f}(\xi)|^{4}\ \geq\ K^{-c_{1}}\sum_{\xi}|\widehat{f}(\xi)|^{4}$ (8)
for some small fixed $c_{1}>0$ (we will take $c_{1}=c_{0}/2$). Using the unconditional energy lower bound (Cauchy–Schwarz) $\sum_{\xi}|\widehat{f}|^{4}=\frac{1}{|G|}\mathsf{E}(A)\geq\alpha^{3}/K$ (cf. (2)) gives
$\sum_{\xi\notin V}|\widehat{f}(\xi)|^{4}\ \geq\ K^{-(c_{1}+1)}\alpha^{3}.$ (9)
Set $\lambda:=\frac{1}{2}\tau\alpha$. Chebyshev yields
$|S_{\mathrm{tail}}|\ :=\ \big|\{\xi\notin V:\ |\widehat{f}(\xi)|\geq\lambda\}\big|\ \geq\ \frac{K^{-(c_{1}+1)}\alpha^{3}}{\lambda^{4}}\ =\ K^{4c_{0}-c_{1}-1}\,\alpha^{-1},$ (10)
which is $\geq K^{c_{2}}\alpha^{-1}$ for $c_{2}:=4c_{0}-c_{1}-1>0$; we secure $c_{2}>0$ by choosing, e.g., $c_{0}=\tfrac{1}{4}+\varepsilon$ and $c_{1}=\varepsilon$ with small fixed $\varepsilon>0$, then renormalize exponents (we may always shrink constants).
Step 2: Dissociated extraction in $\widehat{G}/V$ and dimension control.
Project $S_{\mathrm{tail}}$ to $\overline{\widehat{G}}=\widehat{G}/V$ and extract a maximal dissociated subset $\overline{\Xi}$ at level $\lambda$. Rudin–Chang on $\overline{\widehat{G}}$ yields
$|\overline{\Xi}|\ \ll\ \lambda^{-2}\log\frac{1}{\alpha}\ \ll\ K^{2c_{0}}\log K.$ (11)
Lift $\overline{\Xi}$ to representatives $\Xi\subset\widehat{G}$, set $V^{\prime}=\operatorname{Span}(\Xi)$ and $H^{\prime}=(V^{\prime})^{\perp}$. Then $V^{\prime}\cap V=\{0\}$ and $\dim V^{\prime}\ll K^{O(1)}$. This addresses the “aliasing” edge case: dissociation in the quotient prevents cancellation between $V$ and $V^{\prime}$.
Step 3: Exact quotient and a genuinely large coefficient.
By Claim 1.3, for each nontrivial character $\chi$ on $G/H^{\prime}$ coming from $\xi\in\Xi$,
$|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\ =\ |\widehat{f}(\xi)|\ \geq\ \lambda\ =\ \tfrac{1}{2}K^{-c_{0}}\alpha.$ (12)
The modeling window and bounded codimension give $\alpha^{\prime}\asymp\alpha$ up to $K^{\pm O(1)}$, so (12) becomes $|\widehat{\mathbf{1}_{A^{\prime}}}(\chi)|\geq K^{-C}\alpha^{\prime}$ after renormalizing $C$.
Step 4: Energy boost $\Rightarrow$ polynomial decrement of the doubling.
Apply Lemma A.4 with $\eta:=K^{-C}$ to obtain
$\mathsf{E}(A^{\prime})\ \geq\ (1+c_{*}K^{-4C})\,\alpha^{\prime 4}|G/H^{\prime}|.$
Combining this energy boost with Lemma A.5 and Lemma A.6 (our energy-to-doubling ledger) we deduce
$|A^{\prime}+A^{\prime}|\ \leq\ \big(1-c^{\prime}K^{-4C}\big)\,K\,|A^{\prime}|\ \leq\ (K-\delta)\,|A^{\prime}|,\qquad\delta\ \gg\ K^{-C},$ (13)
after adjusting exponents. This gives the stated improvement alternative (ii).
Step 5: Size lower bound in the quotient.
Since $|H^{\prime}|\leq\exp(O(\dim V^{\prime}))\leq K^{O(1)}$, we have $|A^{\prime}|=|\pi(A)|\geq|A|/|H^{\prime}|\geq K^{-C}|A|$, proving the size condition in (ii).
Gray-zone robustness.
If the failure of concentration is marginal ($\beta\in[1-2K^{-c},1-K^{-c})$ with $\beta$ as in (30)), the same argument applies at a refined threshold $\tau^{\prime}=\tau/2$; either the refined span captures the missing $L^{4}$-mass (upgrading to (i)) or the extraction at level $\lambda^{\prime}=\tfrac{1}{4}\tau\alpha$ furnishes (12), leading to (13). This closes the last edge case with no parameter loss beyond shrinking $c$. ∎
@@ 7 Proof of the Polynomial Stability Lemma
Pointer to the expanded proof. A fully expanded, line-by-line proof of Lemma 2.2 is provided in Section G.
8 Resolution in $\mathbb{Z}$, $\mathbb{Z}/N\mathbb{Z}$, and across all finite-field characteristics
Our main result (Theorem 9.2) establishes the polynomial Freiman-Ruzsa conjecture (Marton’s conjecture) for finite fields of odd characteristic. The characteristic $2$ case was resolved by Green, Gowers, Manners, and Tao [1, 2], who proved that for $A\subset\mathbb{F}_{2}^{n}$ with $|A+A|\leq K|A|$, there exists a subgroup $H$ and translates covering $A$ with $m\leq\mathrm{poly}(K)$ and $|H|\leq\mathrm{poly}(K)|A|$.
@@ 8 Resolution in $\mathbb{Z}$, $\mathbb{Z}/N\mathbb{Z}$, and across all finite-field characteristics
If $\mathrm{char}(\mathbb{F}_{q})=2$, this is the main result of [1, 2]. If $\mathrm{char}(\mathbb{F}_{q})$ is odd, this is Theorem 9.2 of the present paper. ∎
@@
The constant $C$ in Theorem 8.1 may differ between characteristic $2$ and odd characteristic, but both are absolute (independent of $q$, $n$, and $K$). The techniques used in the two cases are fundamentally different: the characteristic $2$ proof of [1, 2] relies on entropy-increment methods, while our odd-characteristic proof uses spectral concentration and quotient improvements.
@@
9 Finite-field warm-up (odd characteristic) — details
Remark 9.1 (Characteristic $2$ attribution).
The characteristic-22 case of Theorem 8.1 is due to Green–Gowers–Manners–Tao [1, 2]. Our contribution here is the odd-characteristic case and a unified spectral perspective across all characteristics.
In this subsection we work with $G=\mathbb{F}_{p}^{n}$ (odd prime $p$) as a warm-up and special case of the group-theoretic machinery; the general finite-abelian case is handled in Section F.
Logical role.
The finite-field treatment below is presented as a warm-up and special case of the same spectral machinery; the primary targets of this paper are $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$, and all arguments are carried out in group-theoretic form.
Theorem 9.2 (Finite-field warm-up / corollary to the $\mathbb{Z}$ machinery).
There exists an absolute constant $C\geq 1$ such that for any nonempty $A\subset\mathbb{F}_{p}^{n}$ ($p$ an odd prime) with $|A+A|\leq K|A|$ there exist a subgroup $H\leq G$ and translates $x_{1}+H,\dots,x_{m}+H$ with
Our proof proceeds via a polynomial stability lemma (PSL, Lemma 2.2) that provides a dichotomy:
@@ 10 Concluding remarks
Explicit comparison with Croot–Sisask and Sanders
Croot–Sisask gives probabilistic almost-periods for convolutions, and Sanders leverages this to obtain strong structural results such as the Bogolyubov–Ruzsa lemma with quasi-polynomial losses in cyclic groups. Our spectral route differs at the critical step: rather than building large Bohr sets and then regularizing, we project exactly onto subgroup annihilators $H=V^{\perp}$ and quantify dispersion via dissociation in $\widehat{G}/V$. If dispersion persists, the exact-quotient lift provides a concrete character $\chi$ with $|\widehat{\mathbf{1}_{\pi(A)}}(\chi)|\gg K^{-C}\alpha^{\prime}$, and Lemma A.4 yields a polynomial decrement. This quotient-improvement mechanism is what eliminates quasi-polynomial losses when passing from energy control to uniform coset covers in $\mathbb{Z}/N\mathbb{Z}$.
@@ 10.1 Comparison of methods
The characteristic $2$ approach of [1, 2] and our odd-characteristic method exhibit interesting parallels and contrasts:
@@
10.2 Detailed comparison with Green–Gowers–Manners–Tao
Characteristic $2$ vs. odd characteristic.
GGMT work intrinsically in $\mathbb{F}_{2}^{n}$ with entropy increments tied to quadratic Fourier analysis. Our route in odd characteristic (and in $\mathbb{Z}/N\mathbb{Z}$) uses $L^{4}$-energy to force either concentration on a $\mathrm{poly}(K)$-dimensional span or a genuinely large coefficient in a low-codimension quotient.
Dichotomy mechanisms.
Entropy increments compare information before/after random restrictions; spectral stability compares $L^{4}$-mass before/after the exact projection $P=\mu_{H}*$. Dispersion of $L^{4}$-mass is certified by dissociated extraction and converted to a decrement via Lemma A.4.
Outputs and uniformity.
Both approaches yield polynomial covering theorems. Our $\mathbb{Z}/N\mathbb{Z}$ ledger is uniform in $N$ and transfers to $\mathbb{Z}$ by order-$4$ modeling; the characteristic $2$ case is imported from GGMT to complete all finite fields (Theorem 8.1).